OpenStudy (anonymous):

use polynomial long division to divide: -4x^3+24x^2-24x+4/ x^2-5x + 1= 1x+1x^2+1x+1= 2x+1x^2+1 Is this correct?

OpenStudy (anonymous):

or would it be\[-4x^3+24x^2-24x+4\div x^2-5x+1= -4x+4 =4-4x \]